OpenStudy (anonymous):

Suppose that x* means 5x2 - x, then what does (y+3)* mean?

OpenStudy (anonymous):

That should be squared..

OpenStudy (anonymous):

No, those are the only instructions. But 5x is squared. Do I maybe plug in (y+3)?

OpenStudy (anonymous):

It was a guess ha

OpenStudy (anonymous):

Maybe 5(y+3)^2-(y+3) 5(y+3) (y+3)* = 5y+15?

OpenStudy (anonymous):

So you only take a y out of the second term? Not the first squared y?

OpenStudy (anonymous):

Yes so you get 5(y+3)(y+3)-(y+3) So the (y+3) crosses out leaving just 5(y+3)?

OpenStudy (anonymous):

I hope so.

Directrix (directrix):

(y+3)* = 5y^2+30y+45 - (y + 3) (y+3)* = 5y^2+30y+45 -1 (y + 3) (y+3)* = 5y^2+30y+45 -y -3 (y+3)* = 5y^2 +29y - 42